- Press Release
- Nov 28, 2022
NASA Space Station On-Orbit Status 15 Mar 2004
All ISS systems continue to function nominally, except as noted previously or below. Underway: Week 21 of Increment 8.
Before breakfast, both crewmembers completed another session of the periodic Russian medical experiment protocols PZEh-MO-7 (calf volume measurement) and PZEh-MO-8 (body mass measurement). FE Alex Kaleri set up the MO-8 “scales” equipment and later broke it down and stowed it away. (Last time done: 3/3). [Calf measurements (left leg only) are taken with the IZOG device, a custom-sewn fabric cuff that fits over the calf, using the knee and lower foot as fixed reference points, to provide a rough index of deconditioning in zero-G and effectiveness of countermeasures. For determining body ass in zero-G, where things are weightless (but not massless), the Russian IM “scales” measure the inertial forces that arise during the oscillatory motion of a mass driven by two helical metering springs with known spring constants. By measuring the time period of each oscillation of the unknown mass (the crewmember) and comparing it to the period of a known mass, the crewmember’s mass is calculated by the computer and displayed. For operational monitoring and data recording the special IFEP (in-flight examination program) software is used on the medical equipment computer (MEC).]
CDR/SO Michael Foale reviewed the upcoming renewed CBOSS-FDI (Cellular Biotechnology Support Systems-Fluid Dynamics Investigation) activities. After the VTR was turned on for recording camcorder video of the ops, Mike worked with the FDI for several hours, practicing syringe injection rates into a TCM (tissue culture module) and experimenting with new methods of improved mixing for future experiments.
Later, the Science Officer activated the MSG (Microgravity Science Glovebox) to make preparations for new PFMI (Pore Formation & Migration Investigation) experiments starting next Thursday (3/18). The video part of the planned PFMI/g-LIMIT activities was deferred. The MSG was later deactivated again.
The FE completed his regular monitoring/servicing of the BIO-5 Rasteniya-2 (“Plants-2”) zero-G plant growth experiment, checking water level in the Lada-4 greenhouse water container and seed status.
The crew continued Elektron troubleshooting by performing leak checks on the Liquid Unit (BZh) separator. No leaks were reported.
To raise ppO2 (oxygen partial pressure) in the cabin atmosphere, another two SFOG candles were burned today. This makes a total of six candles used since last Saturday. None of the six candles experienced any anomalies.
To make room on the timeline for the extensive Elektron troubleshooting, an MO-9/Urolux functionality test, regular maintenance of protective mesh screens of the TsV ventilation fans in the FGB, and the FE’s mandatory 30-min. Crew Medical Officer (CMO) proficiency training were deferred to a later date.
Alex Kaleri worked at the replaceable pump panel (SPN) in the FGB on a performance evaluation of the module’s #2 coolant fluid pump (VGK2-2) of the internal thermal loop #2. [For the test, he changed electrical connectors of the onboard cable network (BKS) at the SPN to switch from VGK2-2 to the primary pump (VGK2-1).]
Mike Foale recharged both batteries of the MedOps defibrillator equipment to full capacity. [Each battery was left in for ~3.5 hrs. and its open-circuit voltage tested with the volt/amp scopemeter, then removed and stowed again. Nominally, the Defib has a battery installed at all times, but with this particular unit the PDIM (power data interface module) would overcharge the batteries.]
Later in the day, Kaleri was scheduled to perform his third round of the Russian MBI-8 “Profilaktika” (countermeasures) fitness test series, on the VELO stationary bike ergometer (last time done: 12/17/03). [There will be two more days for this test series, one with the NS-1 Load Trainer, the other with the TVIS treadmill. Test procedure is identical to the Russian MO-5 assessment, but in addition to the nominal procedure, it calls for the use of the TEEM-100M gas analyzer, measurement of the lactate level in the subject’s blood with the AccuSport device, and subjective evaluation of physical exertion levels during the test. Results are entered on a log sheet. TEEM and ECG (electrocardiograph) data are transferred to the borrowed U.S. SSC6 IBM 760XD laptop, also on a tape cassette (“Cardiocassette-2000”), and prepared for later downlink via Regul-Packet comm. The lactate levels were called down to specialists standing by at TsUP.]
The crew completed the weekly maintenance of the TVIS treadmill. [The machine has worked smoothly after last week’s extensive repair, during which the stabilizer wire ropes were inspected for signs of fraying and found in good shape, not requiring replacement.]
Mike Foale undertook the regular periodic hatch seal inspection on all six USOS (U.S. segment) hatches, viz., Node (forward & aft starboard), Lab (aft), and Joint Airlock, in support of regular ACS (atmospheric control system) maintenance.
Sasha brought the IMS (inventory management system) record up to date to account for the completed Progress 13P cargo transfers, then prepared the daily “delta” file for automatic export/import to update the IMS database. He also performed the daily routine maintenance of the SOZh life support system.
The CDR worked out on TVIS and RED exerciser (Kaleri’s exercise being part of MBI-8).
Since the brief failure of communications between CMG-2 (control moment gyro #2) and its controlling MDM computer early yesterday morning (ref. status report 3/14), the malfunction has occurred two more times until now. Each time, nominal operation of CMG-2 was restored quickly by the automated FDIR (Failure Detection, Identification & Recovery) system power-cycling (rebooting) the MDM. This CMG-2 comm failure has been seen many times before in the past.
Over the weekend, elevated vibrations were again observed in CMG-3 (0.05g instead of nominal 0.02g). The phenomenon continues to be investigated, and there is some suspicion that it could be connected to interaction between CMG-3 and CMG-4. [CMG-1 has been inoperable for a long time.]
Discussions are underway to add a ninth docked day to the Soyuz 8S crew exchange period in late April. [A decision on this change, which would provide the visiting ESA guest cosmonaut André Kuipers additional science time, is expected by end of this week.]
Another OBT (onboard training) exercise on fire emergency procedures is in preparation for tomorrow, with both Mission Control Centers participating.
Planning is underway for a new robotics proficiency activity for the crew with the SSRMS (Space Station Remote Manipulator System) next Thursday (3/18). [Among else, one of the PDGFs (power & data grapple fixtures) will be tested, which Jim Voss and Susan Helms of Expedition 2 had found to be “sticky”.]
Today’s CEO (Crew Earth Observations) targets, limited in XPOP attitude by flight rule constraints on the use of the science window, which is available for only ~1/4 of each orbit when not facing forward (in “ram”), were Darling River floods, Australia (Dynamic event. ISS pass paralleled the course of southeast Australia’s major drainage: a mapping pass along the river floodplain was requested), Angolan Biomass Burning (good pass across the central Angola: looking left and right for early dry-season fires), Etosha dry lake, Namibia (Dynamic event. Crew was asked to document this large dry lake [in northern Namibia’s semidesert] as the wet season ends. Nadir pass over the usually dry, white lake bed), and Patagonian Glaciers (opportunities on the east/Patagonian side of the South Andes ice fields may have existed. Shooting the smaller glacier tongues).
CEO images can be viewed at the websites.
See also the website “Space Station Challenge” at http://voyager.cet.edu/iss/
U.S. and Russian Segment Status (as of today, 2:21pm EST).
Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):
- Elektron O2 generator is Off. Vozdukh CO2 scrubber is On. U.S. CDRA CO2 scrubber is on Standby (ready in dual-bed mode). TCCS (trace contaminant control subsystem) is operating. SM Gas Analyzer has been calibrated and is used for ppO2 and ppCO2 monitoring. MCA (major constituents analyzer) is in Life Extending Mode (LEM). BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode. RS air conditioner SKV-1 is On, SKV-2 is Off (repair now completed; to be tested ASAP).
- SM Working Compartment: Pressure (mmHg) — 742; temperature (deg C) — 26.5; ppO2 (mmHg) — 140.4; ppCO2 (mmHg) — 3.7;
- SM Transfer Compartment: Pressure (mmHg) — 743; temperature (deg C) — 20.1.
- FGB Cabin: Pressure (mmHg) — 744; temperature (deg C) — 24.7.
- Node: Pressure (mmHg) — 740.71; temperature (deg C) — 24.9 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
- U.S. Lab: Pressure (mmHg) — 741.79; temperature (deg C) — 24.9; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
- Joint Airlock (Equip. Lock): Pressure (mmHg) — 741.99; temperature (deg C) — 28.4; shell heater temp (deg C) — 26.9, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
- PMA-1: Shell heater temp (deg C) — 24.0
- PMA-2: Shell heater temp (deg C) — 9.6
(n/a = data not available)
Electrical Power Systems (EPS):
- Both P6 channels fully operational. BGA (beta gimbal assembly) 2B and 4B both in in Autotrack (suntracking) and bias-angled 43 deg. for drag reduction (“sun slicer”)
- SM batteries: Battery #7 is in “Cycle” mode; all other batteries (7) are in “Partial Charge” mode.
- FGB batteries: Battery #6 is off-line (capacity restoration mode, ROM); battery #3 is in “Cycle” mode; all other batteries (4) are in “Partial Charge” mode.
- Plasma Contactor Unit PCU-1 is in Standby mode; PCU-2 is in Standby mode.
Command & Data Handling Systems:
- C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
- GNC-1 MDM is prime; GNC-2 is Backup.
- INT-2 is operating; INT-1 is Off.
- EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
- LA-1, LA-2 and LA-3 MDMs are all operating.
- PL-1 MDM is Off; PL-2 MDM is Operational.
- APS-1 (automated payload switch #1) and APS-2 are both On.
- SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1 dropped out 11/22).
- SM Central Computer (TsVM): 2 redundant lanes (of 3) operational (string #3 dropped out 10/22).
- Total propellant load available: 3948 kg (8704 lb) as of 3/5/04 [SM(755) + FGB(2534) + Progress M(0) + Progress M-1(659)]. (Capability: SM — 860 kg; FGB — 6120 kg).
Attitude Control Systems:
- 3 CMGs on-line (CMG-1 failed).
- State vector source — U.S. SIGI-1 (GPS)
- Attitude source — U.S. SIGI-1 (GPS)
- Angular rate source — RGA-1
XPOP (x-axis perpendicular to orbit plane = “sun-fixed” [yaw: 0.5 deg, pitch: -9.0 deg., roll: 0 deg]), with CMG TA (thruster assist).
Communications & Tracking Systems:
- FGB MDM-1 is powered Off; FGB MDM-2 is operational.
- All other Russian communications & tracking systems are nominal.
- *S-band is operating nominally (on string 2).
- Ku-band is operating nominally.
- Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
- Video subsystem operating nominally.
- HCOR (high-rate communications outage recorder) is operating nominally.
- SSRMS/Canadarm2 based at Lab PDGF/LEE A, powered on both strings.
- MBS: KA power on both strings.
- MT: latched and mated at WS4.
- POA: KA power on both strings.
- RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.
ISS Orbit (as of this morning, 8:19am EST [= epoch]):
- Mean altitude — 368.0 km
- Apogee — 375.1km
- Perigee — 360.8 km
- Period — 91.91 min.
- Inclination (to Equator) — 51.628 deg
- Eccentricity — 0.0010633
- Orbits per 24-hr. day — 15.67
- Mean altitude loss last 24 hours — 80 m
- Revolutions since FGB/Zarya launch (Nov. ’98) — 30362
For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see